OpenStudy (anonymous):

find the derivative of the function y=9cot n(theta)

OpenStudy (anonymous):

\[\sqrt[\csc \tan \left[ \lfloor \ge \supset \rfloor \right]]{?}\]

OpenStudy (anonymous):

Agreed^. Here's my work.

OpenStudy (anonymous):